If Lynn can type a page in p minutes, what piece of the page can she do in 5 minutes?
Answer by josmiceli(19441)

Her rate of typing is
( 1 page ) / ( p minutes )
The formula to use is:
( fraction of a page typed ) = ( rate of typing ) x ( time to type it )
----------------
given:
( fraction of a page typed ) = ( 1/p ) x ( 5 min )
( fraction of a page typed ) = 5/p
She can type 5/p of a page in 5 minutes
----------------
To check this, suppose p = 7
She can type 1 page in 7 minutes
She can type 5/p = 5/7 of a page in 5 minutes